# Fundamental limitations on the recoverability of quantum processes

### v1

@@ Abstract
Quantum information processing and computing tasks can be understood as quantum networks, comprising quantum states and channels and possible physical transformations on them. It is hence pertinent to estimate the change in informational content of quantum processes due to physical transformations they undergo. The physical transformations of quantum states are described by quantum channels, while the transformations of quantum channels are described by quantum superchannels. In this work, we determine fundamental limitations on how well can the physical transformation on quantum channels be undone or reversed, which are of crucial interest to design and benchmark quantum information and computation devices. In particular, we refine the quantum data processing inequality for quantum channels by strengthening the monotonicity inequality of quantum relative entropy of quantum channels under the action of an arbitrary quantum superchannel. We identify a class of quantum superchannels, which appears to be quantum superchannel analogue of quantum subunital channels, under the action of which the entropy of an arbitrary quantum channel is nondecreasing.
@@ I.1 Motivation and background
Quantum theory lies at the core of our understanding of the natural phenomena at the small scale and plays a fundamental role in the rapidly growing area of quantum information processing and computing. Traditionally, quantum states are used as the information carriers and they need to be appropriately manipulated to perform the desired protocols or algorithms, e.g., quantum key distribution ${}^{+}$ (20); DBWH (21) and protocols implying the advantage in quantum computation Sho (94); Gro (96); BGK (18); these manipulations are quantum state transformations. The physical transformations of quantum states of systems are described by completely positive (CP), trace-preserving (TP) maps in quantum theory and are referred to as quantum channels. Achieving desired manipulations of quantum states require precise application of specific quantum channels. A thorough understanding of quantum channels, therefore, is not only essential for the theoretical insights but also pivotal for the realization of quantum computers, sensing and information processing devices, and quantum Internet NC (00); DM (03); Hol (12); ${}^{+}$ (15); Wil (13); SSHD (23).
@@
However, with the rapid increase in our capabilities to handle and manipulate small scale systems, there is a resounding global effort to realize existing quantum technologies in laboratories and invent new ones. This necessitates a convenient and overarching framework to handle quantum states, measurements, and quantum channels succinctly and efficiently in a unified way. Quantum networks comprising quantum states, measurements, and quantum channels provide such a framework, where information can be encoded in quantum channels and/or quantum states depending on the desired tasks ${}^{+}$ (06); Das (19); DW19b ; ${}^{+}$ (23); then the manipulation of information can be treated in a unified way by the higher order transformations of quantum networks. Considering quantum networks as elementary quantum objects, their transformations would then describe the most general kind of physical transformations of quantum information. Examples of transformations of quantum channels include among others the programmable quantum channels (taking quantum state as input and outputting a quantum channel) DP (05); GBW (21) and the cloning of unknown unitaries CDP08c . The transformations of quantum channels are achieved by the action of quantum superchannels–linear transformations that map quantum channels into quantum channels CDP08a ; CDP08b ; Życ (08); Per (17); BP (19). In this work, we shed light on the irreversibility of the transformation of quantum channels under the action of quantum superchannels. In particular, we derive entropic inequalities that provide non-trivial lower bounds on the entropy change and data processing inequality pertaining to the transformation of quantum channels under the action of quantum superchannels.
@@
The entropic functionals attributed to quantum channels, together with their inherent properties, furnishes both qualitative and quantitative methodologies to analyse quantum processes from the perspective of quantum information theory. Entropic inequalities, like uncertainty relations CBTW (17), second laws of thermodynamics ${}^{+}$ (15), data processing inequality Rus (02), etc. have been instrumental in devising quantum communication and computation protocols. The quantum relative entropy $D\left(\cdot\;\middle\|\;\cdot\right)$, a quantum generalization of Kullback-Leibler divergence, is one of the core entropic functions and is indispensable to analysis of distinguishability problems in the quantum information theory. $D\left(\rho\;\middle\|\;\sigma\right)$ between a state $\rho$ of an arbitrary dimensional quantum system, finite or infinite, and a positive semidefinite operator $\sigma$ is defined as Ume (62) [^1] The logarithm can be taken with respect to the base $2$ without loss of generality.
@@
The quantum data processing inequality has several ramifications in the quantum information theory and is one of the widely invoked inequalities to derive limitations in quantum processing and computing tasks Rus (02); Hol (12); ${}^{+}$ (17); DBWH (21); SSC (23); KW (24). A consequence of the above inequality is a second law of thermodynamics like inequality which states that the von Neumann entropy  [^2] We simply refer to the von Neumann entropy functional as the entropy. We also discuss generalized entropy functions. The use of the entropy for the von Neumann entropy will be clear from the context. of a quantum state is nondecreasing under the action of an unital channel AU (77, 82), i.e., $S(\mathcal{N}(\rho))-S(\rho)\geq 0$ for any quantum channel $\mathcal{N}$ that preserves identity operator, i.e., $\mathcal{N}(\mathbbm{1})=\mathbbm{1}$, and arbitrary quantum state $\rho$. It is known that the von Neumann entropy of any quantum state is also nondecreasing for subunital channels AU (77, 82), meaning that $S(\mathcal{N}(\rho))-S(\rho)\geq 0$ for quantum channels $\mathcal{N}$ satisfying $\mathcal{N}(\mathbbm{1})\leq\mathbbm{1}$. The input and output dimensions of unital channels are the same whereas the output dimension of a subunital channel is greater than or equal to the input dimension.
The quantum data processing inequality (Eq. (2)) for states is saturated, i.e., $D\left(\rho\;\middle\|\;\sigma\right)=D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)$, if and only if there exists a completely positive, trace preserving map $\mathcal{P}$, called the Petz recovery map, such that $\mathcal{P}\circ\mathcal{N}(\rho)=\rho$ and $\mathcal{P}\circ\mathcal{N}(\sigma)=\sigma$ Pet (86, 88). This result has found crucial applications in quantum error correction BK (02); MN (12) among others. The quantum data processing inequality has been recently improved to the following. For arbitrary quantum states $\rho$ and $\sigma$, and an arbitrary quantum channel $\mathcal{N}$, there exists a universal recovery map $\mathcal{P}^{\mathrm{R}}_{\sigma,\mathcal{N}}$ (a completely positive and trace nonincreasing map that depends only on $\sigma$ and $\mathcal{N}$ with no dependence on $\rho$) such that ${}^{+}$ (18)
$D\left(\rho\;\middle\|\;\sigma\right)-D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)\geq-\log F\left(\rho,\mathcal{P}^{\mathrm{R}}_{\sigma,\mathcal{N}}\circ\mathcal{N}(\sigma)\right)\geq 0,$ (3)
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}_{1}^{2}$ is the fidelity between $\rho$ and $\sigma$. $\norm{\mathsf{X}}_{1}=\operatorname{tr}{\sqrt{\mathsf{X}^{\ast}\mathsf{X}}}$ is the trace-norm of $\mathsf{X}$. There is also a strengthened (refined) inequality BDW (16) for entropy change of an arbitrary state under the action of an arbitrary positive and trace-preserving map $\mathcal{N}$,
@@
where $\mathcal{N}^{\ast}$ denotes the adjoint map of the linear map $\mathcal{N}$. The above inequality is a strengthened version of entropy change under quantum channels as quantum channels are special cases of positive and trace-preserving maps. Notice that $D\left(\rho\;\middle\|\;\mathcal{N}^{\ast}\circ\mathcal{N}(\rho)\right)\geq 0$ for all subunital channels $\mathcal{N}$ and quantum states $\rho$, thus $\mathcal{N}^{\ast}$ can be deemed as a universal recovery map (channel) for unital channels $\mathcal{N}$ in the context of entropy change.
@@ I.2 Our contribution and main results
We strengthen the data processing inequality for quantum channels by deriving nontrivial lower bound to the monotonicity of relative entropy of quantum channels under the action of a quantum superchannel. That is, the following lower bound holds for the monotonicity of the relative entropy $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ between quantum channels $\mathcal{N}$ and $\mathcal{M}$ with a desirable property under the action of a quantum superchannel $\Theta$,
@@ I.3 Structure of the paper
The structure of this paper is as follows. In Section II, we fix our notations and present the preliminaries and background necessary to derive the main results of this paper. In Section II, we discuss some properties of the space of CP maps, quantum channels, and superchannels; and define the notion of representing maps for superchannels. In Section III, we define the entropy of a quantum channel via an axiomatic approach and we also extend the definition of entropy for quantum channels with infinite-dimensional input of output space. We also prove several bounds on channel entropy gain for a wide class of channels including tele-covariant channels. In Section IV, we introduce the notion of the sufficiency of superchannels and we prove a stronger version of data processing inequality and entropy gain of quantum channels, with a remainder term which depends on recovery supermap. In Section V, we define the generalized divergence and entropy functional for superchannels and we discuss some properties of both the functionals. Finally, in Section VI, we give the concluding remarks of our findings. We also define a divergence functional for the higher order process, we then also provide a necessary condition for the functional to be a generalized divergence. Using this definition of generalized divergence, we define the entropy functional for higher-order processes.
@@ II Preliminaries: Standard notations, definitions, and facts
Let $\mathcal{L}\left(A,B\right)$ be the space of all linear maps from $\mathcal{L}\left({A}\right)$ to $\mathcal{L}\left({B}\right)$. We denote a linear map $\mathcal{N}\in\mathcal{L}\left(A,B\right)$ as $\mathcal{N}_{A\rightarrow B}$. We will use both notations $\mathcal{N}_{A\to B}$ and $\mathcal{N}:\mathcal{L}(A)\to\mathcal{L}(B)$, interchangeably for a linear map $\mathcal{N}$ whenever there is no ambiguity. A linear map $\mathcal{N}_{A\rightarrow B}$ is called positive if it maps elements of $\mathcal{L}\left({A}\right)_{+}$ to the elements of $\mathcal{L}\left({B}\right)_{+}$, and completely positive if $\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow B}$ (where $\mathrm{i}d_{R}$ denotes the identity superoperator acting on a reference system $R$) is a positive map for arbitrary size of the reference system $R$. A map $\mathcal{N}_{A\rightarrow B}$ is called trace preserving if for all $\mathsf{X}\in{\mathcal{L}}({A})$, $\operatorname{tr}\left(\mathcal{N}_{A\rightarrow B}(\mathsf{X})\right)=\operatorname{tr}\left(\mathsf{X}\right)$, and it is called trace nonincreasing if for all $\mathsf{X}\in{\mathcal{L}}(A)$, $\operatorname{tr}\left(\mathcal{N}_{A\rightarrow B}(\mathsf{X})\right)\leq\operatorname{tr}(\mathsf{X})$. A quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({B}\right)$ is a linear, completely positive and trace-preserving (CPTP) map which describes a physically meaningful evolution of a quantum system. A linear map $\mathcal{N}_{A\rightarrow B}$ is called sub-unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)\leq\mathbbm{1}_{{B}}$, unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)=\mathbbm{1}_{{B}}$, and super-unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)>\mathbbm{1}_{{B}}$. The adjoint $\mathcal{N}^{*}_{B\rightarrow A}$ of a linear map $\mathcal{N}_{A\rightarrow B}$ is the unique linear map that satisfies the following equation:
@@
Quantum superchannels.— A linear map $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called a supermap which maps elements of $\mathcal{L}\left(A,B\right)$ to elements of $\mathcal{L}\left(C,D\right)$. A supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called completely CP-preserving if the map $\mathrm{id}\otimes\Theta:\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(C,D\right)$ is CP-preserving for for all choices of the Hilbert spaces $A^{\prime}$ and $B^{\prime}$, where $\mathrm{id}$ denotes the identity supermap. A supermap is called TP-preserving if it maps a trace-preserving map into a trace-preserving map. A quantum superchannel is defined as a completely CP-preserving and TP-preserving supermap. It is clear from the definition that a quantum superchannel maps a quantum channel to a quantum channel.
Let $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{C\rightarrow D}$ be any two quantum channels related via a superchannel $\Theta$ as $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)=\mathcal{M}_{C\rightarrow D}$. The action of $\Theta$ on $\mathcal{N}_{A\rightarrow B}$ can be physically realized as follows CDP08b :
@@
Space of quantum channels and superchannels.–Here we discuss the properties of the space of quantum channels and superchannels. We first introduce a notion of inner product on $\mathcal{L}(A,B)$ as the following lemma.
@@
Representing Map.— Let $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ be a supermap. If this map is completely CP-preserving and TP-preserving then it is called a quantum superchannel. The supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ induces a map $T:\mathcal{L}\left({A\otimes B}\right)\rightarrow\mathcal{L}\left({C\otimes D}\right)$, at the level of Choi operators, and is given by
@@
We have already shown a supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)$ is CP, where $\Lambda_{\Theta}$ can be viewed as a map from $\mathcal{L}(A\otimes C)$ to $\mathcal{L}(B\otimes D)$, and $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ is a basis for $\mathcal{L}(A,B)$. Hence, we can determine its complete positivity by checking the positivity of its Choi operator $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$. For detailed proof, see Appendix B.3. Note that $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ is not trace-preserving, but it preserves the trace only for those $\mathsf{X}\in\mathcal{L}\left(A\otimes{B}\right)$ for which $\Gamma_{\mathsf{X}}$ is trace-preserving. But, we can transform the map $\mathsf{T}$ to a map $\mathsf{T}^{{}^{\prime}}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left(C\otimes{D}\right)$ in such a way that $\mathsf{T}^{{}^{\prime}}$ is trace preserving and $\mathsf{T}(\mathsf{X})=\mathsf{T}^{{}^{\prime}}(\mathsf{X})$, whenever $\Gamma_{\mathsf{X}}$ is trace preserving. The explicit form of the map $\mathsf{T}^{{}^{\prime}}$ is given by the following expression:
$\mathsf{T}^{{}^{\prime}}(\mathsf{X})=\mathsf{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0},$ (20)
where $\sigma_{0}\in\mathcal{L}(C\otimes D)$ is a fixed state. In particular, $\mathsf{T}^{{}^{\prime}}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)=\mathsf{T}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)$ holds for all quantum channels $\mathcal{N}$, if $\mathsf{T}$ does not change the space of operators, or $|A|=|C|$. A given representing map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left(C\otimes{D}\right)$ induces a map $\Lambda:\mathcal{L}(A,B)\rightarrow\mathcal{L}(C,D)$ as follows,

### v2

@@ Abstract
Quantum information processing and computing tasks can be understood as quantum networks, comprising quantum states and channels and possible physical transformations on them. It is hence pertinent to estimate the change in informational content of quantum processes due to physical transformations they undergo. The physical transformations of quantum states are described by quantum channels, while the transformations of quantum channels are described by quantum superchannels. In this work, we determine fundamental limitations on how well the physical transformation on quantum channels can be undone or reversed, which are of crucial interest to design and benchmark quantum information and computation devices. In particular, we refine (strengthen) the quantum data processing inequality for quantum channels under the action of quantum superchannels. We identify a class of quantum superchannels, which appears to be the superchannel analogue of subunital quantum channels, under the action of which the entropy of an arbitrary quantum channel is nondecreasing. We also provide a refined inequality for the entropy change of quantum channels under the action of an arbitrary quantum superchannel.
@@ I.1 Motivation and background
Quantum theory lies at the core of our understanding of the natural phenomena at the small scale and plays a fundamental role in the rapidly growing area of quantum information processing and computing. Traditionally, quantum states are used as the information carriers and they need to be appropriately manipulated to perform the desired protocols or algorithms, e.g., quantum key distribution XMZ+ (20); DBWH (21) and protocols implying the advantage in quantum computation Sho (94); Gro (96); BGK (18); these manipulations are quantum state transformations. The physical transformations of quantum states of systems are described by completely positive (CP), trace preserving (TP) maps in quantum theory and are referred to as quantum channels. Achieving desired manipulations of quantum states require precise application of specific quantum channels. A thorough understanding of quantum channels, therefore, is not only essential for the theoretical insights but also pivotal for the realization of quantum computers, sensing and information processing devices, and quantum Internet NC (00); DM (03); Hol (12); SSD+ (15); Wil (13); SSHD (23).
@@
However, with the rapid increase in our capabilities to handle and manipulate small scale systems, there is a resounding global effort to realize existing quantum technologies in laboratories and invent new ones. This necessitates a convenient and overarching framework to handle quantum states, measurements, and quantum channels succinctly and efficiently in a unified way. Quantum networks comprising quantum states, measurements, and quantum channels provide such a framework, where information can be encoded in quantum channels and/or quantum states depending on the desired tasks SNB+ (06); Das (19); DW19b ; PLY+ (23); then the manipulation of information can be treated in a unified way by the higher order transformations of quantum networks. Considering quantum networks as elementary quantum objects, their transformations would then describe the most general kind of physical transformations of quantum information. Examples of transformations of quantum channels include among others the programmable quantum channels (taking quantum state as input and outputting a quantum channel) DP (05); GBW (21) and the cloning of unknown unitaries CDP08c . The transformations of quantum channels are achieved by the action of quantum superchannels–linear transformations that map quantum channels into quantum channels CDP08a ; CDP08b ; Życ (08); Per (17); BP (19). In this work, we shed light on the irreversibility of the transformation of quantum channels under the action of quantum superchannels. In particular, we derive entropic inequalities that provide non-trivial lower bounds on the entropy change and data processing inequality pertaining to the transformation of quantum channels under the action of quantum superchannels.
@@
The entropic functionals attributed to quantum channels, together with their inherent properties, furnishes both qualitative and quantitative methodologies to analyse quantum processes from the perspective of quantum information theory. Entropic inequalities, like uncertainty relations CBTW (17), second laws of thermodynamics BHN+ (15), data processing inequality Rus (02), etc. have been instrumental in devising quantum communication and computation protocols. The quantum relative entropy $D\left(\cdot\;\middle\|\;\cdot\right)$, a quantum generalization of Kullback-Leibler divergence, is one of the core entropic functions and is indispensable to analysis of distinguishability problems in the quantum information theory. $D\left(\rho\;\middle\|\;\sigma\right)$ between a state $\rho$ of an arbitrary dimensional quantum system, finite or infinite, and a positive semidefinite operator $\sigma$ is defined as Ume (62) [^1] The logarithm can be taken with respect to the base $2$ without loss of generality.
@@
The quantum data processing inequality has several ramifications in the quantum information theory and is one of the widely invoked inequalities to derive limitations in quantum processing and computing tasks Rus (02); Hol (12); BSF+ (17); DBWH (21); SSC (23); KW (24). A consequence of the above inequality is a second law of thermodynamics like inequality which states that the von Neumann entropy  [^2] We simply refer to the von Neumann entropy functional as the entropy. We also discuss generalized entropy functions. The use of the entropy for the von Neumann entropy will be clear from the context. of a quantum state is nondecreasing under the action of a unital channel AU (77, 82), i.e., $S(\mathcal{N}(\rho))-S(\rho)\geq 0$ for any quantum channel $\mathcal{N}$ that preserves identity operator, i.e., $\mathcal{N}(\mathbbm{1})=\mathbbm{1}$, and arbitrary quantum state $\rho$. It is known that the von Neumann entropy of any quantum state is also nondecreasing for subunital channels AU (77, 82), meaning that $S(\mathcal{N}(\rho))-S(\rho)\geq 0$ for quantum channels $\mathcal{N}$ satisfying $\mathcal{N}(\mathbbm{1})\leq\mathbbm{1}$. The input and output dimensions of unital channels are the same whereas the output dimension of a subunital channel is greater than or equal to the input dimension.
The quantum data processing inequality (Eq. (2)) for states is saturated, i.e., $D\left(\rho\;\middle\|\;\sigma\right)=D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)$, if and only if there exists a completely positive, trace preserving map $\mathcal{P}$, called the Petz recovery map, such that $\mathcal{P}\circ\mathcal{N}(\rho)=\rho$ and $\mathcal{P}\circ\mathcal{N}(\sigma)=\sigma$ Pet (86, 88). This result has found crucial applications in quantum error correction BK (02); MN (12) among others. The quantum data processing inequality has been recently improved to the following. For arbitrary quantum states $\rho$ and $\sigma$, and an arbitrary quantum channel $\mathcal{N}$, there exists a universal recovery map $\mathcal{P}^{\mathrm{R}}_{\sigma,\mathcal{N}}$ (a completely positive and trace nonincreasing map that depends only on $\sigma$ and $\mathcal{N}$ with no dependence on $\rho$) such that JRS+ (18)
$D\left(\rho\;\middle\|\;\sigma\right)-D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)\geq-\log F\left(\rho,\mathcal{P}^{\mathrm{R}}_{\sigma,\mathcal{N}}\circ\mathcal{N}(\rho)\right)\geq 0,$ (3)
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}_{1}^{2}$ is the fidelity between $\rho$ and $\sigma$. $\norm{\mathsf{X}}_{1}=\operatorname{tr}{\sqrt{\mathsf{X}^{\ast}\mathsf{X}}}$ is the trace-norm of $\mathsf{X}$. There is also a strengthened (refined) inequality BDW (16) for entropy change of an arbitrary state under the action of an arbitrary positive and trace preserving map $\mathcal{N}$,
@@
where $\mathcal{N}^{\ast}$ denotes the adjoint map of the linear map $\mathcal{N}$. The above inequality is a strengthened version of entropy change under quantum channels as quantum channels are special cases of positive and trace preserving maps. Notice that $D\left(\rho\;\middle\|\;\mathcal{N}^{\ast}\circ\mathcal{N}(\rho)\right)\geq 0$ for all subunital channels $\mathcal{N}$ and quantum states $\rho$, thus $\mathcal{N}^{\ast}$ can be deemed as a universal recovery map (channel) for unital channels $\mathcal{N}$ in the context of entropy change.
@@ I.2 Our contribution and main results
We show that the quantum relative entropy of quantum channels is nonincreasing under the action of a wide class of CP-preserving and TP-preserving supermaps (see Proposition 2). We strengthen the data processing inequality for quantum channels by deriving nontrivial lower bound to the monotonicity of relative entropy of quantum channels under the action of quantum superchannels. That is, the following lower bound holds for the monotonicity of the relative entropy $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ between quantum channels $\mathcal{N}$ and $\mathcal{M}$ with a desirable property under the action of a quantum superchannel $\Theta$,
@@ I.3 Structure of the paper
The structure of this paper is as follows. In Section II, we fix our notations and present the preliminaries and background necessary to derive the main results of this paper. In Section II, we discuss some properties of the space of CP maps, quantum channels, and superchannels; and define the notion of representing maps for superchannels. In Section III, we define the entropy of a quantum channel via an axiomatic approach and we also extend the definition of entropy for quantum channels with infinite-dimensional input of output space. We also prove several bounds on channel entropy gain for a wide class of channels including tele-covariant channels. In Section IV, we introduce the notion of the sufficiency of superchannels and we prove a stronger version of data processing inequality and entropy gain of quantum channels, with a remainder term which depends on recovery supermap. In Section V, we define the generalized divergence and entropy functional for superchannels and we discuss some properties of both the functionals. Finally, in Section VI, we give the concluding remarks of our findings. We also define a divergence functional for the higher order process, and provide a necessary condition for the functional to be a generalized divergence. Using this definition of generalized divergence, we define the entropy functional for higher-order processes.
@@ II Preliminaries: Standard notations, definitions, and facts
Let $\mathcal{L}\left(A,B\right)$ be the space of all linear maps from $\mathcal{L}\left({A}\right)$ to $\mathcal{L}\left({B}\right)$. We denote a linear map $\mathcal{N}\in\mathcal{L}\left(A,B\right)$ as $\mathcal{N}_{A\rightarrow B}$. We will use both notations $\mathcal{N}_{A\to B}$ and $\mathcal{N}:\mathcal{L}(A)\to\mathcal{L}(B)$, interchangeably for a linear map $\mathcal{N}$ whenever there is no ambiguity. A linear map $\mathcal{N}_{A\rightarrow B}$ is called positive if it maps elements of $\mathcal{L}\left({A}\right)_{+}$ to the elements of $\mathcal{L}\left({B}\right)_{+}$, and completely positive if $\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow B}$ (where $\mathrm{i}d_{R}$ denotes the identity superoperator acting on a reference system $R$) is a positive map for arbitrary size of the reference system $R$. A map $\mathcal{N}_{A\rightarrow B}$ is called trace preserving if for all $\mathsf{X}\in{\mathcal{L}}({A})$, $\operatorname{tr}\left(\mathcal{N}_{A\rightarrow B}(\mathsf{X})\right)=\operatorname{tr}\left(\mathsf{X}\right)$, and it is called trace nonincreasing if for all $\mathsf{X}\in{\mathcal{L}}(A)$, $\operatorname{tr}\left(\mathcal{N}_{A\rightarrow B}(\mathsf{X})\right)\leq\operatorname{tr}(\mathsf{X})$. A quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({B}\right)$ is a linear, completely positive and trace preserving (CPTP) map which describes a physically meaningful evolution of a quantum system. A linear map $\mathcal{N}_{A\rightarrow B}$ is called sub-unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)\leq\mathbbm{1}_{{B}}$, unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)=\mathbbm{1}_{{B}}$, and super-unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)>\mathbbm{1}_{{B}}$. The adjoint $\mathcal{N}^{*}_{B\rightarrow A}$ of a linear map $\mathcal{N}_{A\rightarrow B}$ is the unique linear map that satisfies the following equation:
@@
Quantum superchannels.— A linear map $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called a supermap which maps elements of $\mathcal{L}\left(A,B\right)$ to elements of $\mathcal{L}\left(C,D\right)$. A supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called completely CP-preserving if the map $\mathrm{id}\otimes\Theta:\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(C,D\right)$ is CP-preserving for for all choices of the Hilbert spaces $A^{\prime}$ and $B^{\prime}$, where $\mathrm{id}$ denotes the identity supermap. A supermap is called TP-preserving if it maps a trace preserving map into a trace preserving map. A quantum superchannel is defined as a completely CP-preserving and TP-preserving supermap. It is clear from the definition that a quantum superchannel maps a quantum channel to a quantum channel.
Let $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{C\rightarrow D}$ be any two quantum channels related via a superchannel $\Theta$ as $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)=\mathcal{M}_{C\rightarrow D}$. The action of $\Theta$ on $\mathcal{N}_{A\rightarrow B}$ can be physically realized as follows CDP08b ; Gou19a :
@@
Space of quantum channels and superchannels.— Here we discuss the properties of the space of quantum channels and superchannels. We first introduce a notion of inner product on $\mathcal{L}(A,B)$ as the following lemma.
@@
Representing Map.— Let $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ be a supermap. If this map is completely CP-preserving and TP-preserving then it is called a quantum superchannel. The supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ induces a map $\mathsf{T}:\mathcal{L}\left({A\otimes B}\right)\rightarrow\mathcal{L}\left({C\otimes D}\right)$, at the level of Choi operators, and is given by
@@
We have already shown a supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)$ is CP, where $\Lambda_{\Theta}$ can be viewed as a map from $\mathcal{L}(A\otimes C)$ to $\mathcal{L}(B\otimes D)$, and $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ is a basis for $\mathcal{L}(A,B)$. Hence, we can determine its complete positivity by checking the positivity of its Choi operator $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$. For detailed proof, see Appendix B.3.
Note that the representing map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ preserves the trace for those $\mathsf{X}\in\mathcal{L}\left(A\otimes{B}\right)$ for which $\Gamma_{\mathsf{X}}$ is trace preserving but in general $\mathsf{T}$ could neither be a trace nonincreasing map nor a trace increasing map. However, we can transform the map $\mathsf{T}$ into a map $\mathsf{T}{{}^{\prime}}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left(C\otimes{D}\right)$ in such a way that $\mathsf{T}{{}^{\prime}}$ is trace preserving and $\mathsf{T}(\mathsf{X})=\mathsf{T}{{}^{\prime}}(\mathsf{X})$ holds whenever $\Gamma_{\mathsf{X}}$ is trace preserving. The explicit form of the map $\mathsf{T}{{}^{\prime}}$ is given by the following expression:
$\mathsf{T}{{}^{\prime}}(\mathsf{X})=\mathsf{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0},$ (20)
where $\sigma_{0}\in\mathcal{L}(C\otimes D)$. In particular, $\mathsf{T}{{}^{\prime}}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)=\mathsf{T}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)$ holds for all quantum channels $\mathcal{N}$, if $\mathsf{T}$ does not change the space of operators, or $|A|=|C|$. Observe that $\mathsf{T}^{\prime}$ is CP if $\mathsf{T}$ is CP and trace nonincreasing map and we choose $\sigma_{0}$ to be a fixed state. For the case when $\mathsf{T}$ is neither trace nonincreasing map nor trace increasing map, we have derived necessary and sufficient conditions for $\mathsf{T}^{\prime}$ to be a CP map in terms of Choi operator of $\mathsf{T}$ in Appendix B.4. Let us consider a class $\mathsf{SCT}$ of superchannels such that for any superchannel $\Theta\in\mathsf{SCT}$, the representing map $\mathsf{T}$ of $\Theta$ satisfies our necessary and sufficient criteria such that $\mathsf{T}^{\prime}$ is CP (see Appendix B.4). In section IV , we will use this class of superchannels to establish several results for the channel recovery problem.
A given representing map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left(C\otimes{D}\right)$ induces a map $\Lambda:\mathcal{L}(A,B)\rightarrow\mathcal{L}(C,D)$ as follows,